Woodbridge's stock is currently selling for $26.00 a share but is expected to decrease to either $23.40 or increase to $28.60 a share over the next year. The risk-free rate is 3 percent. What is the current value of a 1-year call option with an exercise price of $26? $1.50 $1.64 $1.72 $1.86 $2.02
Answer #1

Answer : $1.64

Risk free rate (r) = 0.03

Strike Price = $ 26

Spot Price (S) = $ 26

If share goes up than Price (S1) = $ 28.60

value of call if share moves up (C1) = 2.6

% change on the up side= (28.60-26)/26 = 10%

r = 1 +% increase in the share price = 1+0.1 = 1.1

If share goes down than price (S2) = $23.40

Price of call if share moves down (C2) = 0

% decrease in the share price =( 23.40-26)/26 = -10%

d = 1 - % decrease in share price = 0.90

Probability of an up move in the share price (p) = (1+r-d)/u-d   = (1+0.03-0.9)/(1.1-0.9) = 0.65

Probability of a downside move in the share price (1-p) = 0.35

Call Price = [(p*C1)+(1-p) *C2]/(1+r)

C = (0.65*2.6+0.35*0)/1.03

C = 1.64

Therefore the value of call option of $ 26 strike price will be $ 1.64.